# Go Beyond Earth: Understanding Human Actions and Scenes in Microgravity Environments

### v3

@@ 5.1 Benchmark Protocol
Data Split. We partition the dataset into training, validation, and test subsets in a 7:1:2 ratio. Row‐level distribution is as follows: Training set: 9,266 records (69.93% of 13,251); Validation set: 1,329 records (10.03% of 13,251); Test set: 2,656 records (20.04% of 13,251). Video‐level distribution is as follows: Training set: 3,331 videos (69.99% of 4,759); Validation set: 475 videos (9.98% of 4,759); Test set: 953 videos (20.03% of 4,759).
Baselines for Fine-Grained HAR in Microgravity Scenarios. For fine-grained action recognition, we evaluate well-established baselines from video-based human action recognition, including transformer-based models (MViTv1 Fan et al. (2021), MViTv2 Li et al. (2022)) and CNN-based models (I3D Carreira & Zisserman (2017), SlowFast Feichtenhofer et al. (2019), X3D Feichtenhofer (2020), C2D Feichtenhofer et al. (2019)). These widely used architectures cover both paradigms and allow us to assess generalization to the unique spatiotemporal dynamics of microgravity. All models are initialized with Kinetics400 Kay et al. (2017) pretrained weights for better convergence.
Baseline methods for Video Captioning and QA in Microgravity Scenarios. For video captioning and question answering in microgravity, we evaluate strong baselines, including open-source models (VideoChatGPT Li et al. (2023), mPLUG-Owl-3 Ye et al. (2025), LLaVA-Next Li et al. (2024), VideoLLaVa Lin et al. (2024), Qwen-2.5-VL Bai et al. (2025), InternVideo Wang et al. (2022)) and closed-source models (GPT-4o Hurst et al. (2024), Gemini 1.5 Pro Reid et al. (2024)). Open-source models offer reproducibility via public weights and code, while closed-source models serve as upper-bound references. This mix enables both transparent analysis and comprehensive evaluation of video-language understanding in microgravity.
@@
Evaluation Metrics for Fine-Grained HAR. Our evaluation metrics include mAP@0.5, F1 score, recall, and AUROC, all calculated using the macro method. Among these, mAP@0.5 is the primary metric for measuring average detection accuracy per category and thus comprehensively evaluating the model’s action recognition performance in a microgravity environment. Per-class threshold sweeps are provided in the appendix.
@@ 5.2 Result Analysis for Fine-Grained Human Action Recognition
Note: All models have been pretrained on Kinetics400 Kay et al. (2017) dataset and continually trained on MicroG-4M. TC denotes the temporal configuration (frame length × sampling rate). #Params indicates the number of parameters (in millions, M).
@@
Model Test Result Arch TC Backbone Pretrain Fine-tune mAP (%) F1-score (%) Recall (%) AUROC (%) Slow Feichtenhofer et al. (2019) 8x8 R50 He et al. (2016) Kinetics 400 Kay et al. (2017) AVA v2.2 Gu et al. (2018) $16.24$ $2.67$ $1.99$ $73.83$ SlowFast Feichtenhofer et al. (2019) 32x2 R101 He et al. (2016) Kinetics 600 Carreira et al. (2018) AVA v2.2 Gu et al. (2018) $23.81$ $6.32$ $6.62$ $77.83$
@@
Figure 3: Qualitative results for fine-grained human action recognition in microgravity, where GT denotes ground truth, MicroG-4M indicates predictions from Slow fine-tuned on MicroG-4M, and AVA denotes predictions from the same model fine-tuned on AVA. The MicroG-4M model provides more accurate predictions than its Earth-trained counterpart.
@@ 5.3 Evaluation of Video Captioning Models
The results in Table 5.3 reveal how video-language models perform on the MicroG-4M benchmark, highlighting key challenges introduced by microgravity-specific content. Lexical metrics such as CIDEr and BLEU-4 show low overall values, especially among open-source models, suggesting a significant distributional shift between MicroG-4M and typical pretraining data. The dataset’s domain-specific vocabulary, visually compositional scenes, and semantically dense annotations likely reduce surface-level overlap, which these metrics are sensitive to. In contrast, semantic similarity metrics such as S-BERT and BERTScore remain relatively higher and more consistent, indicating that several models capture the underlying intent even without lexical alignment. This underscores the semantic richness of MicroG-4M, where alternative phrasings and scientific terminology often convey similar meanings. Performance differences further reveal that input frame density and pretraining modality play key roles. InternVideo Wang et al. (2022), which processes $90$ frames sampled within a $3$s window, consistently outperforms other open-source models. This suggests that dense sampling, coupled with video-specific pretraining, enhances the model’s ability to capture subtle spatial patterns and object-scene relationships—features that are particularly important in microgravity scenarios, where visual cues are often atypical or physically ambiguous. Closed-source models, i.e., GPT-4o Hurst et al. (2024) and Gemini 1.5 Pro Reid et al. (2024) achieve better scores on both lexical and semantic metrics, likely due to broader data exposure, larger capacity, or more advanced cross-modal fusion strategies. However, their relatively small performance gains further validate the challenge posed by MicroG-4M. In general, these findings position MicroG-4M as a demanding benchmark for evaluating multimodal models under domain change, highlighting the need for robust spatial reasoning, domain adaptation, and semantically aware generation strategies in unconventional environments.
@@ 5.4 Evaluation of Visual Question Answering Models
Table 5: Experiments of Visual Question Answering (VQA) models on the MicroG-4M benchmark. “#f” denotes the number of input frames used during model inference.
@@
The results in Table 5.4 demonstrate that MicroG-4M presents distinct challenges for visual question answering. Notably, there is a significant divergence between lexical and semantic evaluation metrics, particularly among open-source models. For example, Qwen2.5-VL Bai et al. (2025) yields a BLEU-4 of only 0.65 and a CIDEr score of 2.99, yet achieves the highest S-VQA score in its category (40.65). This contrast suggests that MicroG-4M questions often admit multiple semantically valid answers that differ lexically, such as paraphrased actions, scientific terms, or object references adapted to microgravity settings. This characteristic does not reflect inconsistency in evaluation, but rather underscores the conceptual and linguistic diversity embedded in the dataset. In addition, the moderate absolute scores of even the top-performing closed-source models, such as GPT-4o Hurst et al. (2024) (CIDEr 33.98, S-VQA 44.56), reveal the difficulty of reasoning over visual content in microgravity. Unlike conventional VQA datasets, MicroG-4M includes visually ambiguous cues, e.g., floating objects, unusual body orientations, and tool manipulations under microgravity that challenge models trained primarily on terrestrial data. This suggests that current pretraining corpora lack sufficient coverage of such scenarios, and that purely scaling model capacity is insufficient for reliable generalization. Interestingly, increasing the number of input frames within the fixed $3$s window does not consistently yield better performance. For example, Gemini 1.5 Pro Reid et al. (2024) processes 16 frames but performs worse than GPT-4o Hurst et al. (2024), which uses only 6 frames. This indicates that dense frame sampling alone is insufficient. Instead, performance depends more critically on the model’s ability to extract semantically salient cues, e.g., astronaut posture, object manipulation, and spatial configurations, from visually subtle or low-motion segments. In microgravity environments, where conventional motion dynamics and object affordances are altered, effective spatial reasoning and cross-modal alignment appear to be more decisive than temporal redundancy. In summary, MicroG-4M reveals key limitations of current VQA systems in addressing domain-specific challenges, particularly those involving spatial complexity, ambiguous motion, and semantically flexible queries inherent to microgravity. Its comprehensive and specialized design establishes it as a valuable testbed for probing the robustness, adaptability, and generalization capabilities of multimodal models well beyond the scope of conventional Earth-based benchmarks.
@@ 6 Conclusion
Acknowledgement
The project served to prepare the SFB 1574 Circular Factory for the Perpetual Product (project ID: 471687386), approved by the German Research Foundation (DFG, German Research Foundation). This work was supported in part by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026). This project is also supported in part by the National Natural Science Foundation of China under Grant No. 62473139, in part by the Hunan Provincial Research and Development Project (Grant No. 2025QK3019), and in part by the Open Research Project of the State Key Laboratory of Industrial Control Technology, China (Grant No. ICT2025B20).
@@ Acknowledgement
Appendix A Technical Limitations
@@
Appendix B Potential Social Impacts
@@
Appendix C Safety and Ethical Discussion
@@
Appendix D Additional Qualitative Analysis
D.1 Fine-Grained Human Action Recognition
We provide qualitative comparisons using representative video samples to illustrate the fine-grained action recognition performance of different models in microgravity scenarios (Figure 4).
The AVA fine-tuned model frequently misinterprets microgravity-specific postures and contexts, exhibiting errors such as misclassifying hatch-crossing as bend/bow or generating unrealistic predictions like detecting a “smoke” action. In contrast, the MicroG-4M fine-tuned models demonstrate improved accuracy by correctly identifying nuanced and compound actions specific to microgravity environments, as illustrated by the more precise recognition of actions like carry/hold, lift/pickup, and put on/off clothing. The highest-performing MicroG-4M model (I3D Non-local) further reduces both false positives and missed detections, highlighting its superior capability in recognizing complex and subtle astronaut activities.
These examples underscore the necessity for specialized training data, such as MicroG-4M, to effectively capture and recognize fine-grained human actions unique to microgravity conditions.
@@ D.2 Video Captioning
D.3 Video Question Answering
@@
Appendix E Data Format Conversion
@@ E.1 Dataset Metadata
Video naming follows the “[source identifier]_[sequence number]” format, where the identifier is the YouTube ID or film name, and the sequence number indicates its temporal position. This structured metadata approach ensures efficient integration between frame-level and video-level annotations, effectively supporting downstream vision-language research in microgravity environments. For comprehensive documentation of the dataset metadata, please refer to the resources provided on our GitHub and Hugging Face repositories.
Appendix F Supplementary for Human Action Recognition (HAR) Task
F.1 Dataset Partitioning
@@
Train 9,266 (69.93%) 3,331 (69.99%)
Val 1,329 (10.03%) 475 (9.98%)
Test 2,656 (20.04%) 953 (20.03%)
Total 13,251 (100.00%) 4,759 (100.00%)
F.2 Dataset Statistics
@@
F.3 Per-Class AP Results
Figures 7 and 8 display the per-class average precision (AP) performance of models fine-tuned on MicroG-4M and AVA datasets, respectively, and evaluated on the MicroG-4M test set. Fine-tuning on MicroG-4M generally achieves higher per-class AP scores, highlighting improved model adaptation to microgravity-specific action patterns. In contrast, AVA fine-tuned models exhibit significant performance degradation, particularly on fine-grained or microgravity-specific actions such as ‘operate spaceship’, ‘float’, or posture-related actions (‘bend/bow’). These differences underscore the critical need for specialized fine-tuning datasets such as MicroG-4M to address the unique challenges posed by microgravity environments in action recognition tasks.
Figure 7: Per-Class AP after Fine-Tuning on MicroG-4M, evaluated on the MicroG-4M test set. All models shown in this figure correspond to those listed in Table 1 of the main text.
Figure 8: Per-Class AP after Fine-Tuning on AVA, evaluated on the MicroG-4M test set. All models shown in this figure correspond to those listed in Table 2 of the main text.
F.4 Cross-Dataset VQA Density and Design Rationale

### v4

@@ 5.1 Benchmark Protocol
Data Split. We partition the dataset into training, validation, and test subsets in a 7:1:2 ratio. Row‐level distribution is as follows: Training set: 9,2739{,}273 records ($69.93\%$ of 13,26113{,}261); Validation set: $1,330$ records ($10.03\%$ of 13,26113{,}261); Test set: 2,6582{,}658 records ($20.04\%$ of 13,26113{,}261). Video‐level distribution is as follows: Training set: 3,3313{,}331 videos ($69.99\%$ of 4,7594{,}759); Validation set: $475$ videos ($9.98\%$ of 4,7594{,}759); Test set: 953 videos ($20.03\%$ of 4,7594{,}759).
Baselines for Fine-Grained HAR. For fine-grained action recognition, we evaluate well-established baselines from video-based human action recognition, including transformer-based models (MViTv1 Fan et al. (2021), MViTv2 Li et al. (2022)) and CNN-based models (I3D Carreira and Zisserman (2017), SlowFast Feichtenhofer et al. (2019), X3D Feichtenhofer (2020), C2D Feichtenhofer et al. (2019)). These widely used architectures cover both paradigms and allow us to assess generalization to the unique spatiotemporal dynamics of microgravity. All models are initialized with Kinetics400 Kay et al. (2017) pretrained weights for better convergence. HAR experiments were conducted on NVIDIA RTX 2080 Ti GPUs and an AMD Ryzen Threadripper 2990WX CPU.
Baseline methods for Video Captioning and QA. For video captioning and question answering in microgravity, we evaluate strong baselines, including open-source models (VideoChatGPT Li et al. (2023), mPLUG-Owl-3 Ye et al. (2025), LLaVA-Next Li et al. (2024), VideoLLaVa Lin et al. (2024), Qwen-2.5-VL Bai et al. (2025), InternVideo Wang et al. (2022)) and closed-source models (GPT-4o Hurst et al. (2024), Gemini 1.5 Pro Reid et al. (2024), Gemini 2.5 Pro Comanici et al. (2025)). Open-source models offer reproducibility via public weights and code, while closed-source models serve as upper-bound references. This mix enables both transparent analysis and comprehensive evaluation of video-language understanding in microgravity. All evaluations were conducted using NVIDIA A100-40 GPUs and Intel Xeon Platinum 8368 CPUs.
@@
Evaluation Metrics for Fine-Grained HAR. Our evaluation metrics include mAP@0.5, F1 score, recall, and AUROC, all calculated using the macro method. Among these, mAP@0.5 is the primary metric for measuring average detection accuracy per category and thus comprehensively evaluating the model’s action recognition performance in a microgravity environment. Per-class threshold sweeps provided in Appendix.
@@ 5.2 Result Analysis for Fine-Grained Human Action Recognition
Note: All models have been pretrained on Kinetics400 dataset Kay et al. (2017) and continually trained on MicroG-4M. TC denotes the temporal configuration (frame length × sampling rate). #Params indicates the number of parameters (in millions, M).
@@
Model Test Result Arch TC Backbone Pretrain Fine-tune mAP (%) F1-score (%) Recall (%) AUROC (%) Slow Feichtenhofer et al. (2019) 4x16 R50 He et al. (2016) Kinetics 400 Kay et al. (2017) AVA v2.2 Gu et al. (2018) $15.74$ $4.15$ $3.35$ $72.62$ Slow Feichtenhofer et al. (2019) 8x8 R50 He et al. (2016) Kinetics 400 Kay et al. (2017) AVA v2.2 Gu et al. (2018) $16.24$ $2.67$ $1.99$ $73.83$ Slow Feichtenhofer et al. (2019) 32x2 R50 He et al. (2016) Kinetics 400 Kay et al. (2017) AVA v2.2 Gu et al. (2018) $11.02$ $0.87$ $0.58$ $67.26$ SlowFast Feichtenhofer et al. (2019) 8x8 R101 He et al. (2016) Kinetics 600 Carreira et al. (2018) AVA v2.2 Gu et al. (2018) $14.95$ $1.99$ $1.35$ $71.01$ SlowFast Feichtenhofer et al. (2019) 32x2 R101 He et al. (2016) Kinetics 600 Carreira et al. (2018) AVA v2.2 Gu et al. (2018) $23.81$ $6.32$ $6.62$ $77.83$ SlowFast Feichtenhofer et al. (2019) 16x8 R101 He et al. (2016) Kinetics 600 Carreira et al. (2018) AVA v2.2 Gu et al. (2018) $15.08$ $0.70$ $0.42$ $69.14$
@@
Figure 4: Qualitative results for fine-grained action recognition in microgravity. GT denotes ground truth; MicroG-4M and AVA refer to models fine-tuned on MicroG-4M and AVA, respectively. The MicroG-4M model yields more accurate predictions.
@@ 5.3 Evaluation of Video Captioning Models
Gemini 2.5 Pro Comanici et al. (2025) 3 3.27 3.10 17.92 13.08 62.82 86.95
The results in Table 5.3 reveal how video-language models perform on the MicroG-4M benchmark, highlighting key challenges introduced by microgravity-specific content. Lexical metrics such as CIDEr and BLEU-4 show low overall values, especially among open-source models, suggesting a significant distributional shift between MicroG-4M and typical pretraining data. The dataset’s domain-specific vocabulary, visually compositional scenes, and semantically dense annotations likely reduce surface-level overlap, which these metrics are sensitive to. In contrast, semantic similarity metrics such as S-BERT and BERTScore remain relatively higher and more consistent, indicating that several models capture the underlying intent even without lexical alignment. This underscores the semantic richness of MicroG-4M, where alternative phrasings and scientific terminology often convey similar meanings. Performance differences further reveal that input frame density and pretraining modality play key roles. InternVideo Wang et al. (2022), which processes $90$ frames sampled within a $3$s window, consistently outperforms other open-source models. This suggests that dense sampling, coupled with video-specific pretraining, enhances the model’s ability to capture subtle spatial patterns and object-scene relationships—features that are particularly important in microgravity scenarios, where visual cues are often atypical or physically ambiguous. Closed-source models, i.e., GPT-4o Hurst et al. (2024), Gemini 1.5 Pro Reid et al. (2024) and Gemini 2.5 Pro Comanici et al. (2025) achieve better scores on both lexical and semantic metrics, likely due to broader data exposure, larger capacity, or more advanced cross-modal fusion strategies. However, their relatively small performance gains further validate the challenge posed by MicroG-4M. In general, these findings position MicroG-4M as a demanding benchmark for evaluating multimodal models under domain change, highlighting the need for robust spatial reasoning, domain adaptation, and semantically aware generation strategies in unconventional environments.
@@ 5.4 Evaluation of Visual Question Answering Models
Table 5: Experiments of Visual Question Answering (VQA) models on the MicroG-4M benchmark.
@@
The results in Table 5.4 demonstrate that MicroG-4M presents distinct challenges for visual question answering. Notably, there is a significant divergence between lexical and semantic evaluation metrics, visible in both open- and closed-source models. For example, Qwen2.5-VL Bai et al. (2025) yields a BLEU-4 of only 0.65 and a CIDEr score of 2.99, yet achieves the highest S-VQA score in its category (40.65). A similar pattern appears for Gemini 2.5 Pro Comanici et al. (2025), which attains the highest S-VQA score among closed-source models (44.75) while exhibiting very low CIDEr and BLEU-4 scores, largely because it tends to answer simple questions with long, explanatory responses. This contrast suggests that MicroG-4M questions often admit multiple semantically valid answers that differ lexically, such as paraphrased actions, scientific terms, or object references adapted to microgravity settings. This characteristic does not reflect inconsistency in evaluation, but rather underscores the conceptual and linguistic diversity embedded in the dataset. In addition, the moderate absolute scores of even the top-performing closed-source models, such as GPT-4o Hurst et al. (2024) (CIDEr 33.98, S-VQA 44.56), reveal the difficulty of reasoning over visual content in microgravity. Unlike conventional VQA datasets, MicroG-4M includes visually ambiguous cues, e.g., floating objects, unusual body orientations, and tool manipulations under microgravity that challenge models trained primarily on terrestrial data. This suggests that current pretraining corpora lack sufficient coverage of such scenarios, and that purely scaling model capacity is insufficient for reliable generalization. Interestingly, increasing the number of input frames within the fixed $3$s window does not consistently yield better performance. For example, Gemini 1.5 Pro Reid et al. (2024) processes 16 frames but performs worse than GPT-4o Hurst et al. (2024), which uses only 6 frames, and Gemini 2.5 Pro Comanici et al. (2025) achieves the best S-VQA score with just 3 frames. This indicates that dense frame sampling alone is insufficient. Instead, performance depends more critically on the model’s ability to extract semantically salient cues, e.g., astronaut posture, object manipulation, and spatial configurations, from visually subtle or low-motion segments. In microgravity environments, where conventional motion dynamics and object affordances are altered, effective spatial reasoning and cross-modal alignment appear to be more decisive than temporal redundancy. In summary, MicroG-4M reveals key limitations of current VQA systems in addressing domain-specific challenges, particularly those involving spatial complexity, ambiguous motion, and semantically flexible queries inherent to microgravity. Its comprehensive and specialized design establishes it as a valuable testbed for probing the robustness, adaptability, and generalization capabilities of multimodal models well beyond the scope of conventional Earth-based benchmarks.
@@ 6 Conclusion
Acknowledgments
The project is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB-1574 – 471687386. This work was supported in part by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026). This work is also supported in part by the National Natural Science Foundation of China under Grant No. 62473139, in part by the Hunan Provincial Research and Development Project (Grant No. 2025QK3019), and in part by the State Key Laboratory of Autonomous Intelligent Unmanned Systems (the opening project number ZZKF2025-2-10). This research was partially funded by the Ministry of Education and Science of Bulgaria (support for INSAIT, part of the Bulgarian National Roadmap for Research Infrastructure).
Ethics Statement
This work constructs a benchmark from publicly available videos released by official space agencies and publicly distributed films. No private recordings, surveillance footage, or restricted-access materials are included. For public release, we distribute annotations and metadata only, and require users to obtain source videos independently through legal channels. This avoids redistribution of copyrighted audiovisual content and supports compliance with institutional and regional regulations.
The dataset depicts professional activities of publicly visible astronauts and actors in occupational settings. While publicly disseminated material typically does not require additional individual consent for research use, downstream users are instructed to follow their institutions’ IRB/ethics guidelines and applicable privacy regulations. No personally identifying information beyond what is inherently visible in the source videos is collected or released, and metadata exclude personal identifiers.
The dataset does not include satellite imagery or sensitive geolocation data. Film clips are screened for physical plausibility and safety consistency, and implausible scenes are removed. For vision–language components, large language models are used to generate and rank candidate questions under interrogative diversity, with human verification of final questions and answers. Unanswerable “Not mentioned” cases are retained to discourage hallucination and overconfident speculation.
To mitigate misuse, the resource is released for research purposes under terms that prohibit re-identification attempts, surveillance applications, or violations of platform policies and content owners’ rights. We will honor takedown requests from rightsholders and provide a mechanism to remove or amend entries if concerns arise. Overall, the release policy combines metadata-only distribution and clear usage terms, which aims to balance openness and reproducibility with privacy, safety, and copyright considerations.
@@ Acknowledgments
Appendix A The Use of Large Language Models (LLMs)
We use LLMs in two limited ways. During dataset construction, an LLM proposes candidate VQA questions per 3 s clip; humans filter, verify, and finalize both questions and answers, and captions are human written. For benchmarking, several LLM-based systems are evaluated as baselines under a unified inference setup. LLMs do not assign action labels or alter evaluation protocols. We also used an LLM for language polishing of the manuscript, and all technical content and conclusions were written and verified by the authors.
Appendix B Technical Limitations
@@
Appendix C Potential Social Impacts
@@
Appendix D Safety and Ethical Discussion
@@
Appendix E Additional Qualitative Analysis
E.1 Fine-Grained Human Action Recognition
We provide qualitative comparisons using representative video samples to illustrate the fine-grained action recognition performance of different models in microgravity scenarios (Figure 5). These examples reveal two consistent signatures that drive the domain gap between Earth-trained and microgravity-tuned models.
First, gravity-dependent person-movement classes are fragile. When the vertical reference is ambiguous or decoupled from posture, the AVA-finetuned model projects microgravity poses onto terrestrial categories such as bend/bow, crouch/kneel, sit, or stand. Representative cases include: hatch crossing labeled as bend/bow (upper panel, column 1); free-floating or foot-restrained standing either mislabeled as sit (upper panel, column 3; lower panel, column 1) or not detected at all (upper panel, column 4; lower panel, column 5); sideways crouch/kneel missed entirely (lower panel, last column); and microgravity sit mapped to lie/sleep due to orientation mismatch (upper panel, column 5). These observations indicate that person-movement classes that implicitly assume an upright axis are unstable in orbit.
Second, transient force direction and short-contact object manipulation are difficult. Short episodes of lift/pick up and put down are detected inconsistently or absorbed into background motion (upper panel, column 2; lower panel, column 2). The action hit (an object) receives low confidence and is suppressed by the detection threshold, yielding no action label (lower panel, rightmost column). In contrast, device-centric actions with clear visual anchors and sustained contact, most notably carry/hold (an object), are comparatively stable and predicted consistently across models in the same examples (upper panel, column 2; lower panel, columns 2–4).
@@ E.2 Video Captioning
E.3 Video Question Answering
@@
Appendix F Data Format Conversion
@@ F.1 Dataset Metadata
Video naming follows the "[source identifier]_[sequence number]" format, where the identifier is the YouTube ID or film name, and the sequence number indicates its temporal position. This structured metadata approach ensures efficient integration between frame-level and video-level annotations, effectively supporting downstream vision-language research in microgravity environments. For comprehensive documentation of the dataset metadata, please refer to the resources provided on our GitHub and Hugging Face repositories.
Appendix G Supplementary for Human Action Recognition (HAR) Task
G.1 Dataset Partitioning
@@
Train 9,273 (69.93%) 3,331 (69.99%)
Val 1,330 (10.03%) 475 (9.98%)
Test 2,658 (20.04%) 953 (20.03%)
Total 13,261 (100.00%) 4,759 (100.00%)
G.2 Dataset Statistics
@@
G.3 Per-Class AP Results
Building on the aggregate results in Table 1 and Table 2, we analyze per-class AP on the MicroG-4M test set. Figures 814 report the per-class curves with the 50 AVA-derived MicroG-4M actions ordered as in Table 9.
Paired comparison.
[…]
Across remaining backbones.
For C2D, I3D, MViT family, and X3D, MicroG-4M fine-tuning generally improves the posture-sensitive actions and short-contact manipulations above, yet several classes remain hard across architectures: pull (an object), push (an object), close (e.g., a door, a box), take a photo, text on/look at a cellphone, fight/hit (a person), and take (an object) from (a person). These actions share low visual salience or momentary cues such as small handheld objects, brief hand–hand transfers, global camera/scene drift in spacecraft cabins, which limits per-class AP even after in-domain training (see Figures 12, 13,  14, and 15).
Backbone-specific observations.
MViTv1, MViTv2, and X3D exhibit comparatively limited gains after MicroG-4M fine-tuning. Two factors are consistent with this trend. First, temporal windowing and sampling may not balance long drift with short contact, which requires both extended context and reliable short-term discrimination. Second, inductive biases differ: backbones with stronger local encoding and non-local augmentation align better with microgravity-specific postures and contact cues than global-attention–dominant designs with weaker local bias or very lightweight CNNs without non-local pathways. These patterns are visible in the per-class plots for X3D and MViT (Figures 14 and 15), and they are consistent with the aggregate results in Tables 1 and 2, which together explain the test mAP plateau and indicate that in-domain fine-tuning recovers part of the gap while several action categories remain challenging in orbit.
Figure 8: Per-class AP (%) on the MicroG-4M test set for Slow Feichtenhofer et al. (2019). Training regime: MicroG-4M fine-tuning. AP@0.5 over the 50 MicroG-4M action classes; class order matches Table 9. Legend uses $L\times S$ to denote clip length $\times$ sampling stride; settings match Table 1.
Figure 9: Per-class AP (%) on the MicroG-4M test set for Slow Feichtenhofer et al. (2019). Training regime: AVA fine-tuning with zero-shot evaluation on MicroG-4M. AP@0.5 over the 50 AVA-derived action classes. Legend uses $L\times S$ to denote clip length $\times$ sampling stride; settings match Table 1.
Figure 10: Per-class AP (%) on the MicroG-4M test set for SlowFast Feichtenhofer et al. (2019). Training regime: MicroG-4M fine-tuning. AP@0.5 over the 50 MicroG-4M action classes; class order matches Table 9. Legend uses $L\times S$ to denote clip length $\times$ sampling stride; settings match Table 1.
Figure 11: Per-class AP (%) on the MicroG-4M test set for SlowFast Feichtenhofer et al. (2019). Training regime: AVA fine-tuning with zero-shot evaluation on MicroG-4M. AP@0.5 over the 50 AVA-derived action classes. Legend uses $L\times S$ to denote clip length $\times$ sampling stride; settings match Table 1.
Figure 12: Per-class AP (%) on the MicroG-4M test set for C2D Feichtenhofer et al. (2019). NLN denotes the non-local augmentation of the backbone. Training regime: MicroG-4M fine-tuning. AP@0.5 over the 50 MicroG-4M action classes; class order matches Table 9. Legend uses $L\times S$ to denote clip length $\times$ sampling stride; settings match Table 1.
Figure 13: Per-class AP (%) on the MicroG-4M test set for I3D Carreira and Zisserman (2017). NLN denotes the non-local augmentation of the backbone. Training regime: MicroG-4M fine-tuning. AP@0.5 over the 50 MicroG-4M action classes; class order matches Table 9. Legend uses $L\times S$ to denote clip length $\times$ sampling stride; settings match Table 1.
Figure 14: Per-class AP (%) on the MicroG-4M test set for MViTv1 Fan et al. (2021) and MViTv2 Li et al. (2022). Training regime: MicroG-4M fine-tuning. AP@0.5 over the 50 MicroG-4M action classes; class order matches Table 9. Legend uses $L\times S$ to denote clip length $\times$ sampling stride; settings match Table 1.
Figure 15: Per-class AP (%) on the MicroG-4M test set for X3D Feichtenhofer (2020). Training regime: MicroG-4M fine-tuning. AP@0.5 over the 50 MicroG-4M action classes; class order matches Table 9. Legend uses $L\times S$ to denote clip length $\times$ sampling stride; settings match Table 1.
G.4 Cross-Dataset VQA Density and Design Rationale